# The Honeycomb Conjecture in normed planes and an alpha-convex variant of a theorem of Dowker

### v1

@@ 3. Results about polygonal norms
The goal of this section to investigate the $\alpha$-honeycomb properties of poygonal norms, and the $\alpha$-Dowker properties of polygons.
@@ 3.1. An algorithm to check the $\alpha$-honeycomb property of a polygonal norm
In this subsection we give a simple algorithm that check if the isoperimetrix of a polygonal norm satisfies the $\alpha$-Dowker, or weak $\alpha$-Dowker property for some given $\alpha\in[0,\infty)$. We assume that $\alpha>0$, and note that our argument can be modified in a natural way for the log-Dowker or weak log-Dowker property.
@@
for values $4\leq n\leq 2k$. Similarly, to check if $M_{iso}$ satisfies the weak $\alpha$-Dowker property, it is enough to check if the inequality in (1.2) is satisfied for $m=3,4,5$ and $6<n\leq 2k$.
We also recall that there is an algorithm, given by Aggarwal, Chang and Yap [1] that computes, for any given convex polygon $P$ with $s$ sides, the area of a minimum area convex $t$-gon for any given $3\leq t<s$, in $O(s^{2}\log s\log t)$ steps.
@@ 2. Step 2
    : Using the algorithm in [1], we compute the quantities $A_{M{iso}}(n)$ for all $3\leq n\leq 2k$. This step requires $\mathcal{O}(k^{3}\log^{2}k)$ steps.
@@ 3. Step 3
    : We can check inequality in (1.2) for $m=3,4,5$ and all $6<n\leq 2k$, in $\mathcal{O}(k)$ steps.
@@ 3.2. Norms where the unit disk is a regular polygon
[…]
@@
Let $3\leq n\leq m$. Let $P$ be a convex $m$-gon, and let $Q$ denote a minimum area convex $n$-gon circumscribed about $P$. Then the midpoint of every side of $Q$ belongs to $P$. Furthermore, if $n\geq 4$, then there is a minimum area convex $m$-gon $Q^{\prime}$ such that there is at most one side of $Q^{\prime}$ that does not contain a side of $P_{n}$, and if $S$ is such a side of $Q^{\prime}$, then the sum of the angles of $Q^{\prime}$ on $S$ is strictly less than $\pi$.
@@
Now, let $Q^{\prime}$ be a minimum area convex $m$-gon circumscribed about $P$ with the property that it has the smallest number of sides that contain only a vertex of $P$. Let $S$ be such a side of $Q^{\prime}$. Then, by the previous paragraph, the vertex $p$ of $P$ on $S$ is the midpoint of $S$. Let $\alpha$ and $\beta$ denote the two angles of $Q^{\prime}$ on $S$. If $\alpha+\beta>\pi$, then slightly rotating the sideline of $Q^{\prime}$ through $S$ about $p$ decreases the area of $Q^{\prime}$, contradicting the assumption that $Q^{\prime}$ is a minimum area circumscribed $m$-gon. If $\alpha+\beta=\pi$, then the same rotation does not change the area of $Q^{\prime}$, and thus, rotating this line until it first reaches another vertex of $S$ we obtain a minimum area circumscribed polygon having strictly less sides that contain only a vertex of $P$. Thus, we have $\alpha+\beta<\pi$. But any convex $n$-gon with $n\geq 4$ has at most one side satisfying this property. ∎
@@
We note that if $P$ is $o$-symmetric and $n$ is even, then the method of the proof of Lemma 3.5, combined with Theorem 2.7, yields that there is a minimum area convex $m$-gon $Q$ circumscribed about $P$ such that $Q$ is $o$-symmetric, and every side of $Q$ contains a side of $P$.
@@
Let $k\geq 2$ and $3\leq n\leq 2k$ be positive integers, and let $2k=nq+r$, where $q,r$ are nonnegative integers, and $r<n$. If $m\geq 4$, then
@@
(3.3) $A(k,3)=\begin{cases}3\tan\frac{\pi}{3},&\text{if}~{}2k=3q,\\ \frac{2}{\cos^{2}\left(\frac{\pi}{2k}\right)}\left(2\sin\left(\frac{q\pi}{k}\right)+\sin\left(\frac{(q+1)\pi}{k}\right)\right),&\text{if}~{}n=3q+1,\\ \frac{2}{\cos^{2}\left(\frac{\pi}{2k}\right)}\left(\sin\left(\frac{q\pi}{k}\right)+2\sin\left(\frac{(q+1)\pi}{k}\right)\right),&\text{if}~{}n=3q+2.\end{cases}$
@@
Subcase 1.1: every side of $Q$ contains a side of $P_{k}$. Without loss of generality, we may assume that the indices of the sides of $P_{k}$ contained in a side of $Q$ are of the form $s_{1},s_{1}+s_{2},\ldots,s_{1}+k_{s}+\ldots+s_{n}=2k$ for some positive integers $s_{i}$, where $s_{i}<k$ for every value of $i$. Since $P_{k}$ is circumscribed about $\mathbf{B}^{2}$, we have
@@
Subcase 1.2, there is a unique side of $Q$ that contains only a vertex of $P_{k}$. Observe that by our assumptions, in this case $n\geq 5$ is odd. From this, by Lemma 3.5, it follows that $k\geq n\geq 5$. Let the sides of $Q$ be denoted by $E_{1},E_{2},\ldots,E_{n}$ in counterclockwise order. For $i=2,\ldots,n$, let $2\alpha_{i}$ denote the turning angle of $Q$ at the common vertex of $E_{i-1}$ and $E_{i}$, and let $2\alpha_{1}$ denote the turning angle at the common vertex of $E_{n}$ and $E_{1}$. Without loss of generality, let $E_{1}$ be the side of $Q$ that contains only a vertex of $P_{k}$. Then, for any $i$ with $3\leq i\leq m$, $\alpha_{i}$ is an integer multiple of $\frac{\pi}{2k}$, and $\sum_{i=1}^{n}\alpha_{i}=\pi$. Since $P_{k}$ is circumscribed about $\mathbf{B}^{2}$, the distance of the line of $E_{1}$ from $o$ is strictly greater than $1$. Let us translate the line of $E_{1}$ towards $o$ until its distance is $1$, and let $Q^{\prime}$ denote the convex $m$-gon obtained in this way (see Figure 2). Then
@@
We intend to construct a convex polygon circumscribed about $P_{k}$ with area smaller than $\operatorname{area}(Q^{\prime})$. To do it, we assume that $s=1$; for the case $s\geq 2$ a straightforward modification of our argument can be given. First, note that as $E_{1}$ contains no side of $P_{n}$, we have that $\alpha_{2}-\alpha_{1}\neq\frac{\pi}{2k}$. Consider the case that $\alpha_{2}-\alpha_{1}>\frac{\pi}{2k}$, then there are integer multiples $\alpha_{2}^{\prime}\geq\alpha_{1}^{\prime}$ of $\frac{\pi}{2k}$ such that $\alpha_{1}^{\prime}+\alpha_{2}^{\prime}=\alpha_{1}+\alpha_{2}$. Then, by the convexity of the $x\mapsto\tan x$ function the expression $\tan\alpha_{1}^{\prime}+\tan\alpha_{2}^{\prime}+\sum_{i=_{3}}^{n}\tan\alpha_{i}$ is the area of a convex $n$-gon $Q^{\prime\prime}$ circumscribed about $P_{k}$, and $\operatorname{area}(Q^{\prime\prime})<\operatorname{area}(Q^{\prime})<\operatorname{area}(Q)$, contradicting our assumption the minimality of the area of $Q$.
@@
Without loss of generality, let $\alpha_{3}>\frac{\pi}{2k}$. Since $\alpha_{3}$ is an integer multiple of $\frac{\pi}{2k}$, this yields that $\alpha_{3}\geq\frac{\pi}{k}>\alpha_{1}$. Set $\alpha_{2}^{\prime}=\alpha_{2}+\frac{\pi}{2k}$, $\alpha_{3}^{\prime}=\alpha_{3}-\frac{\pi}{2k}$. Then, by the convexity of the ${}^{x}\mapsto\tan x$ function, we have that $\sum_{i=1}^{n}\tan\alpha_{i}>\alpha_{1}+\alpha_{2}^{\prime}+\alpha_{3}^{\prime}+\sum_{i=3}^{n}\tan\alpha_{i}$. Here $\alpha_{2}^{\prime}-\alpha_{1}>\frac{\pi}{2k}$, and applying the argument in the previous case obtain a convex $n$-gon $Q^{\prime\prime}$ circumscribed about $P_{k}$ such that $\operatorname{area}(Q^{\prime\prime})<\operatorname{area}(Q^{\prime})$, a contradiction.
@@
Subcase 2.1, every side of $Q$ contains only a vertex of $P_{k}$. Let these vertices of $P_{k}$ be $p_{i},p_{j},p_{l}$, where we can assume that $1\leq i<j<l\leq 2k$. Since these vertices are the midpoints of the corresponding sides of $Q$, $\operatorname{conv}\{p_{i},p_{j},p_{k}\}$ is the midtriangle of $Q$ and hence, the sides of this triangle are parallel to the sides of $Q$. Since $Q$ is circumscribed about $P_{k}$, it follows that $j-i,k-j$ and $i+n-k$ differ by at most one. Thus, by elementary computation we obtain that if $Q$ is a minimal area triangle circumscribed about $P_{k}$, and $\operatorname{bd}(Q)$ contains no side of $T$, then, with the notation $2k=3q+r$, where $q,r$ are nonnegative integers and $r\leq 2$,
@@
Subcase 2.2, exactly one side of $Q$ contains a side of $P_{k}$. Let the sides of $Q_{3}$ be $E_{1},E_{2},E_{3}$, and assume that $[p_{n},p_{1}]\subset E_{1}$, and $p_{i}\in E_{2}$ and $p_{j}\in E_{3}$. Since the latter two vertices are midpoints of $E_{2}$ and $E_{3}$, respectively, we have that $[p_{i},p_{j}]$ is parallel to $[p_{n},p_{1}]$, implying that $j=n+1-i$. Let $q$ denote the midpoint of $E_{1}$, and recall that $q\in[p_{n},p_{1}]$. Note that $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of $Q$, and thus, the area of $Q$ does not depend on the position of $q$. This yields that if $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of a triangle circumscribed about $P_{k}$, then so is $\operatorname{conv}\{p_{1},p_{i},p_{j}\}$ or $\operatorname{conv}\{p_{n},p_{i},p_{j}\}$. Thus, we may apply the argument from Subcase 2.1 and obtain the same quantities for $\operatorname{area}(Q)$ as in (3.4).
Subcase 2.3, exactly two sides of $Q$ contain a side of $P_{k}$. We apply the approach of the previous cases, and lt $\operatorname{conv}\{p,q,r\}$ be the midtriangle of $Q$, where $p$ is a vertex of $P_{k}$. Then, like in Subcase 2.2, we can observe that both $[p,q]$ and $[p,r]$ are parallel to some sides of $P_{k}$. But this yields that both $q,r$ are vertices of $P_{k}$, and we obtain the estimate in (3.4).
Subcase 2.4, every side of $Q$ contains a side of $P_{k}$. Let $T=\operatorname{conv}\{p,q,r\}$ denote the midtriangle of $Q$. Since every side of $T$ is parallel to a side of $P_{k}$, the axial symmetry of $P_{k}$ with respect to the bisector of its every side and a simple geometric observation shows that $p,q,r$ are the midpoints of some sides of $P_{k}$. But then $T$ is axially symmetric with repect to the bisector of its every side, implying that $T$ is regular. But from this we immediately obtain that $2k$ is divisible by $3$, and that $\operatorname{area}(Q)=3\tan\frac{\pi}{3}$.
@@
We need to determine which are the values of $k\geq 2$ satisfy the inequality
(3.5) $\frac{n-6}{n-m}\sqrt{A(k,m)}+\frac{6-m}{n-m}\sqrt{A(k,n)}\geq 2\sqrt{A(k,6)}$
@@
Recall that by definition, $\mathbf{B}^{2}\subset P_{k}$ for all values of $k$, implying that $A_{n}(BB^{2})\leq A_{n}(P_{k})$ for all values of $n$. An elementary computation shows that $A_{n}(\mathbf{B}^{2})=n\tan\frac{\pi}{n}$ for all $n$. Thus, we have
@@
$\sqrt{A(2k,6)}\leq\frac{1}{2}\sqrt{5\tan\frac{\pi}{5}}+\frac{1}{2}\sqrt{7\tan\frac{\pi}{7}}$
for any $q\geq 3$ and $60\leq r<$ 6. Thus, (3.5) holds for any $k\geq 9$. We also observe that since $A(k,n)=\operatorname{area}(P_{k})$ for any $n\geq 2k$, (3.5) clearly holds if $k=2$ or $k=3$. Thus, we are left with the cases $k=4,5,6,7,8$, for which a direct computation of the wo sides of (3.5) for any $3\leq m<6<n\leq 2k$, using a Maple 18.00 software, yields the assertion. ∎

### v2

@@ 3. Results about polygonal norms
The goal of this section is to investigate the $\alpha$-honeycomb properties of polygonal norms, and the $\alpha$-Dowker properties of polygons.
@@ 3.1. An algorithm to check the $\alpha$-honeycomb property of a polygonal norm
In this subsection we give a simple algorithm that checks if the isoperimetrix of a polygonal norm satisfies the $\alpha$-Dowker, or weak $\alpha$-Dowker property for some given $\alpha\in[0,\infty)$. We assume that $\alpha>0$, and note that our argument can be modified in a natural way for the log-Dowker or weak log-Dowker property.
@@
holds for all values $4\leq n\leq 2k$. Similarly, to check if $M_{iso}$ satisfies the weak $\alpha$-Dowker property, it is enough to check if the inequality in (1.2) is satisfied for $m=3,4,5$ and $6<n\leq 2k$.
We also recall that there is an algorithm, given by Aggarwal, Chang and Yap [1] that computes, for any given convex polygon $P$ with $s$ sides, the area of a minimum area circumscribed convex $t$-gon for any given $3\leq t<s$, in $\mathcal{O}(s^{2}\log s\log t)$ steps.
@@ 2. Step 2
    : Using the algorithm in [1], we compute the quantities $A_{M_{iso}}(n)$ for all $3\leq n\leq 2k$. This step requires $\mathcal{O}(k^{3}\log^{2}k)$ steps.
@@ 3. Step 3
    : We check the inequality in (1.2) for $m=3,4,5$ and all $6<n\leq 2k$, in $\mathcal{O}(k)$ steps.
@@ 3.2. Norms where the unit disk is a regular polygon
Note that if the unit disk $M$ of $\mathcal{M}$ is a regular $(2k)$-gon, then $M_{iso}$ is also a regular $(2k)$-gon. Thus, Theorem 3.2 readily follows from combining Theorems 3.3 and 1.9.
@@
Let $3\leq n\leq m$. Let $P$ be a convex $m$-gon, and let $Q$ denote a minimum area convex $n$-gon circumscribed about $P$. Then the midpoint of every side of $Q$ belongs to $P$. Furthermore, if $n\geq 4$, then there is a minimum area convex $n$-gon $Q^{\prime}$ such that there is at most one side of $Q^{\prime}$ that does not contain a side of $P$, and if $S$ is such a side of $Q^{\prime}$, then the sum of the angles of $Q^{\prime}$ on $S$ is strictly less than $\pi$.
@@
Now, let $Q^{\prime}$ be a minimum area convex $n$-gon circumscribed about $P$ with the property that it has the smallest number of sides that contain only a vertex of $P$. Let $S$ be such a side of $Q^{\prime}$. Then, by the previous paragraph, the vertex $p$ of $P$ on $S$ is the midpoint of $S$. Let $\alpha$ and $\beta$ denote the two angles of $Q^{\prime}$ on $S$. If $\alpha+\beta>\pi$, then slightly rotating the sideline of $Q^{\prime}$ through $S$ about $p$ decreases the area of $Q^{\prime}$, contradicting the assumption that $Q^{\prime}$ is a minimum area circumscribed $n$-gon. If $\alpha+\beta=\pi$, then the same rotation does not change the area of $Q^{\prime}$, and thus, rotating this line until it first reaches another vertex of $P$ we obtain a minimum area circumscribed polygon having strictly less sides than $Q^{\prime}$ that contain only a vertex of $P$. Thus, we have $\alpha+\beta<\pi$. But any convex $n$-gon with $n\geq 4$ has at most one side satisfying this property. ∎
@@
We note that if $P$ is $o$-symmetric and $n$ is even, then the method of the proof of Lemma 3.5, combined with Theorem 2.7, yields that there is a minimum area convex $n$-gon $Q$ circumscribed about $P$ such that $Q$ is $o$-symmetric, and every side of $Q$ contains a side of $P$.
@@
Let $k\geq 2$ and $3\leq n\leq 2k$ be positive integers, and let $2k=nq+r$, where $q,r$ are nonnegative integers, and $r<n$. If $n\geq 4$, then
@@
(3.3) $A(k,3)=\begin{cases}3\tan\frac{\pi}{3},&\text{if}~{}2k=3q,\\ \frac{2}{\cos^{2}\left(\frac{\pi}{2k}\right)}\left(2\sin\left(\frac{q\pi}{k}\right)+\sin\left(\frac{(q+1)\pi}{k}\right)\right),&\text{if}~{}2k=3q+1,\\ \frac{2}{\cos^{2}\left(\frac{\pi}{2k}\right)}\left(\sin\left(\frac{q\pi}{k}\right)+2\sin\left(\frac{(q+1)\pi}{k}\right)\right),&\text{if}~{}2k=3q+2.\end{cases}$
@@
Subcase 1.1: every side of $Q$ contains a side of $P_{k}$. Without loss of generality, we may assume that the indices of the sides of $P_{k}$ contained in a side of $Q$ are of the form $s_{1},s_{1}+s_{2},\ldots,s_{1}+s_{2}+\ldots+s_{n}=2k$ for some positive integers $s_{i}$, where $s_{i}<k$ for every value of $i$. Since $P_{k}$ is circumscribed about $\mathbf{B}^{2}$, we have
@@
Subcase 1.2, there is a unique side of $Q$ that contains only a vertex of $P_{k}$. Observe that by our assumptions, in this case $n\geq 5$ is odd. From this, by Lemma 3.5, it follows that $k\geq n\geq 5$. Let the sides of $Q$ be denoted by $E_{1},E_{2},\ldots,E_{n}$ in counterclockwise order. For $i=2,\ldots,n$, let $2\alpha_{i}$ denote the turning angle of $Q$ at the common vertex of $E_{i-1}$ and $E_{i}$, and let $2\alpha_{1}$ denote the turning angle at the common vertex of $E_{n}$ and $E_{1}$. Without loss of generality, let $E_{1}$ be the side of $Q$ that contains only a vertex of $P_{k}$. Then, for any $i$ with $3\leq i\leq n$, $\alpha_{i}$ is an integer multiple of $\frac{\pi}{2k}$, and $\sum_{i=1}^{n}\alpha_{i}=\pi$. Since $P_{k}$ is circumscribed about $\mathbf{B}^{2}$, the distance of the line of $E_{1}$ from $o$ is strictly greater than $1$. Let us translate the line of $E_{1}$ towards $o$ until its distance is $1$, and let $Q^{\prime}$ denote the convex $n$-gon obtained in this way (see Figure 2). Then
@@
We intend to construct a convex polygon circumscribed about $P_{k}$ with area smaller than $\operatorname{area}(Q^{\prime})$. To do it, we assume that $s=1$; for the case $s\geq 2$ a straightforward modification of our argument can be given. First, note that as $E_{1}$ contains no side of $P_{k}$, we have that $\alpha_{2}-\alpha_{1}\neq\frac{\pi}{2k}$. Consider the case that $\alpha_{2}-\alpha_{1}>\frac{\pi}{2k}$, then there are integer multiples $\alpha_{2}^{\prime}\geq\alpha_{1}^{\prime}$ of $\frac{\pi}{2k}$ such that $\alpha_{1}^{\prime}+\alpha_{2}^{\prime}=\alpha_{1}+\alpha_{2}$. Then, by the convexity of the $x\mapsto\tan x$ function the expression $\tan\alpha_{1}^{\prime}+\tan\alpha_{2}^{\prime}+\sum_{i=_{3}}^{n}\tan\alpha_{i}$ is the area of a convex $n$-gon $Q^{\prime\prime}$ circumscribed about $P_{k}$, and $\operatorname{area}(Q^{\prime\prime})<\operatorname{area}(Q^{\prime})<\operatorname{area}(Q)$, contradicting our assumption about the minimality of the area of $Q$.
@@
Without loss of generality, let $\alpha_{3}>\frac{\pi}{2k}$. Since $\alpha_{3}$ is an integer multiple of $\frac{\pi}{2k}$, this yields that $\alpha_{3}\geq\frac{\pi}{k}$. Set $\alpha_{2}^{\prime}=\alpha_{2}+\frac{\pi}{2k}$, $\alpha_{3}^{\prime}=\alpha_{3}-\frac{\pi}{2k}$. Then, by the convexity of the $x\mapsto\tan x$ function, we have that $\sum_{i=1}^{n}\tan\alpha_{i}>\alpha_{1}+\alpha_{2}^{\prime}+\alpha_{3}^{\prime}+\sum_{i=4}^{n}\tan\alpha_{i}$. Here $\alpha_{2}^{\prime}-\alpha_{1}>\frac{\pi}{2k}$, and applying the argument in the previous case we obtain a convex $n$-gon $Q^{\prime\prime}$ circumscribed about $P_{k}$ such that $\operatorname{area}(Q^{\prime\prime})<\operatorname{area}(Q^{\prime})$, a contradiction.
@@
Subcase 2.1, every side of $Q$ contains only a vertex of $P_{k}$. Let these vertices of $P_{k}$ be $p_{i},p_{j},p_{l}$, where we can assume that $1\leq i<j<l\leq 2k$. Since these vertices are the midpoints of the corresponding sides of $Q$, $\operatorname{conv}\{p_{i},p_{j},p_{l}\}$ is the midtriangle of $Q$ and hence, the sides of this triangle are parallel to the sides of $Q$. Since $Q$ is circumscribed about $P_{k}$, it follows that $j-i,l-j$ and $i+2k-l$ differ by at most one. Thus, by an elementary computation we obtain that if $Q$ is a minimal area triangle circumscribed about $P_{k}$, and $\operatorname{bd}(Q)$ contains no side of $T$, then, with the notation $2k=3q+r$, where $q,r$ are nonnegative integers and $r\leq 2$,
@@
Subcase 2.2, exactly one side of $Q$ contains a side of $P_{k}$. Let the sides of $Q$ be $E_{1},E_{2},E_{3}$, and assume that $[p_{2k},p_{1}]\subset E_{1}$, and $p_{i}\in E_{2}$ and $p_{j}\in E_{3}$. Since the latter two vertices are midpoints of $E_{2}$ and $E_{3}$, respectively, we have that $[p_{i},p_{j}]$ is parallel to $[p_{2k},p_{1}]$, implying that $j=2k+1-i$. Let $q$ denote the midpoint of $E_{1}$, and recall that $q\in[p_{2k},p_{1}]$. Note that $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of $Q$, and thus, the area of $Q$ does not depend on the position of $q$. This yields that if $\operatorname{conv}\{q,p_{i},p_{j}\}$ is the midtriangle of a triangle circumscribed about $P_{k}$, then so is $\operatorname{conv}\{p_{1},p_{i},p_{j}\}$ or $\operatorname{conv}\{p_{2k},p_{i},p_{j}\}$. Thus, we may apply the argument from Subcase 2.1 and obtain the same quantities for $\operatorname{area}(Q)$ as in (3.4).
Subcase 2.3, exactly two sides of $Q$ contain a side of $P_{k}$. We apply the approach of the previous cases, and let $\operatorname{conv}\{p,q,r\}$ be the midtriangle of $Q$, where $p$ is a vertex of $P_{k}$. Then, like in Subcase 2.2, we can observe that both $[p,q]$ and $[p,r]$ are parallel to some sides of $P_{k}$. But this yields that both $q,r$ are vertices of $P_{k}$, and we obtain the estimate in (3.4).
Subcase 2.4, every side of $Q$ contains a side of $P_{k}$. Let $T=\operatorname{conv}\{p,q,r\}$ denote the midtriangle of $Q$. Since every side of $T$ is parallel to a side of $P_{k}$, the axial symmetry of $P_{k}$ with respect to the bisector of its every side and a simple geometric observation shows that $p,q,r$ are the midpoints of some sides of $P_{k}$. But then $T$ is axially symmetric with respect to the bisector of its every side, implying that $T$ is regular. From this we immediately obtain that $2k$ is divisible by $3$, and that $\operatorname{area}(Q)=3\tan\frac{\pi}{3}$.
@@
We need to determine which values of $k\geq 2$ satisfy the inequality
(3.5) $\frac{n-6}{n-m}\sqrt{A(k,m)}+\frac{6-m}{n-m}\sqrt{A(k,n)}\geq\sqrt{A(k,6)}$
@@
Recall that by definition, $\mathbf{B}^{2}\subset P_{k}$ for all values of $k$, implying that $A_{n}(\mathbf{B}^{2})\leq A_{n}(P_{k})$ for all values of $n$. An elementary computation shows that $A_{n}(\mathbf{B}^{2})=n\tan\frac{\pi}{n}$ for all $n$. Thus, for any $3\leq m<6<n$, we have
@@
$\sqrt{A(k,6)}\leq\frac{1}{2}\sqrt{5\tan\frac{\pi}{5}}+\frac{1}{2}\sqrt{7\tan\frac{\pi}{7}}$
for any $q\geq 3$ and $60\leq r<$ 6. Thus, (3.5) holds for any $k\geq 9$. We also observe that since $A(k,n)=\operatorname{area}(P_{k})$ for any $n\geq 2k$, (3.5) clearly holds if $k=2$ or $k=3$. Thus, we are left with the cases $k=4,5,6,7,8$, for which a direct computation of the two sides of (3.5) for any $3\leq m<6<n\leq 2k$, using a Maple 18.00 software, yields the assertion. ∎